OpenStudy (anonymous):

I Will Medal And Fan! I need explanation AND answer! Add. 43 + (−44) = −34 + (−39) =

OpenStudy (anonymous):

This could be done simply on your calculator. Just add these, and you get the answer xD Or are you wondering how you get those answers?

OpenStudy (anonymous):

idk I could do that so easily

OpenStudy (anonymous):

43 + (−44) = 43 - 44 = -1

OpenStudy (mrnood):

you have mistyped it -that is not an addition - there should not be an = in the middle. Also - on a calculator you need to be careful of the difference between - (opposite of subtraction) and - (negative number)

OpenStudy (anonymous):

but you can't do -34 + (-39)

OpenStudy (anonymous):

with a calculator

OpenStudy (anonymous):

Yes you can

OpenStudy (mrnood):

are these 2 separate questions?

OpenStudy (anonymous):

nvm. and yes they r

OpenStudy (anonymous):

Yeah are these two separate questions, or are they equal to each other?

OpenStudy (anonymous):

i forgot the -

OpenStudy (anonymous):

ty,

OpenStudy (mrnood):

it is obvious they are not equal to each other -1 not equal -73

OpenStudy (mrnood):

@flatouthero that is not correct at all

OpenStudy (anonymous):

OOPS TYPO negative plus negative is NEGATIVE

OpenStudy (anonymous):

typo

OpenStudy (anonymous):

yeah sry bout that lol xD